A diverging lens with a focal length of -15 cm is placed 16 cm to the right of a converging lens with a focal length of 18 cmA bright object and a viewing screen are separated by a distance of 68.4 cm. Part A At what location(s) between the object anPlease HELP!!! I will rate good for good answers!
Answer #1

distance nce a 33cm Yren &a= -15cm, object -) 4,2 33 cm f = 18 un L2 1 formula, By lens € 1-1 ܩܟ 11 <!- -(-) 18 33 a) 1 ta HEM2 = (39.59 - 16 cm ) u :-(23:59 am) right of leus 1 2 115=tissa) 5) - 23.59 1 -15 V₂= a. 169 cm And Idi= =) Negative sign sh1 1 68.4-d + 1 n 11 d 5 d +6814-d (684-d) (d) 11 -> (68.4-d)d (11)( 68 8.4) =) 68.4d-d2 = 952-4 - d² - 68.4d +752.4=0 d = 68.